OpenStudy (anonymous):

use an inequality to solve: one side of a rectangle is 4 times the other,and the perimeter is not to exceed 150 inches. find the possible values for x, the length of the shorter side.

OpenStudy (anonymous):

Howdy

OpenStudy (anonymous):

lol hi

OpenStudy (anonymous):

tx 4 the fan

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

sooo : lets call one side x

OpenStudy (anonymous):

Now what would be her Loooonger brother called ?

OpenStudy (anonymous):

umm not sure on that one yet

OpenStudy (anonymous):

I am "prodding" you to a mental effort (so to speak)

OpenStudy (anonymous):

"four time longer" the data says - so what is four times x ?

OpenStudy (anonymous):

oh really? lol i didnt know it's been a long time since i been in school

OpenStudy (anonymous):

4x

OpenStudy (anonymous):

im not sure how to work this one

OpenStudy (anonymous):

x + 4x = 150

OpenStudy (anonymous):

sorry twice that 2*(x+4x) = 150

OpenStudy (anonymous):

Can you move forward by urslf. from here ?

OpenStudy (anonymous):

if i understood it better

OpenStudy (anonymous):

i have never had algebra before

OpenStudy (anonymous):

please, please help

OpenStudy (anonymous):

ok so hr. gz. : 2(5x) \[\le 150\] from which follows ( like in Equation ) that 10x les-or-eq 150 so that x l.eq. 15 here is ur answer

OpenStudy (anonymous):

Did it become clearer ?

OpenStudy (anonymous):

im going to practice that one see if i can do it and understand it .. thank you i will leave you alone now lol

OpenStudy (anonymous):

Read on INEQUALITIES (whats the similar and the different from equalities)

OpenStudy (anonymous):

ok thanks i will

OpenStudy (anonymous):

im doing a quiz so i hope there right lol